# Auto-Calibration and Biconvex Compressive Sensing with Applications to Parallel MRI

### v1

@@ 4.4 Dual Certificate
Regarding the exact dual certificate, let $\textbf{V}=\textbf{A}_{S}(\textbf{A}_{S}^{*}\textbf{A}_{S})^{-1}\text{sgn}(\textbf{X}_{0})$, we have $\boldsymbol{\mathcal{Y}}=\textbf{A}^{*}\textbf{A}_{S}(\textbf{A}_{S}^{*}\textbf{A}_{S})^{-1}\text{sgn}(\mathbf{X}_{0})$ is the exact dual certificate. The next lemma deals with bounding the parameter $\tau$ in $\|\textbf{V}\|_{F}\leq\tau\sqrt{s}$ in 3. The bound obtained in the following lemma does not depend on the randomness of H or $sgn(\textbf{X}_{0})$.
@@
[…]
@@
Remark: The condition on $L$ is automatically met due to the bound $L\sim O(kn\log(k^{2}n^{2}N))$ from Lemma (5).
Hence, $\max\limits_{j\in[|S^{c}|]}\|\mathcal{P}_{S^{c}_{j}}\boldsymbol{\mathcal{Y}}\|_{F}<1/2$ with probability at least $1-N^{-\alpha}$ if
@@
5 Simulation
In the first part of the simulations, we demonstrate the requirements of $L$ and $C$ for exact recovery in terms of different values of $k$ and $n$. We calculate the empirical probability of success for different values of $k$ and $n$ and show that the cutoffs for $L$ and $C$ are consistent with the analytically derived bounds.
@@ 5.1 Empirical Success rate for the proposed method
For the first set of experiments, we fix the number of measurements $L=128$ and the underlying signal dimension $N=256$, and solve the optimization problem 17 for varying $(k,n),k,n\in[15]$ and $C\in\{1,2,4\}$. We follow the same setup as in the paper [36] and calculate the empirical probability of success of the proposed method. For each $C$ and $(k,n)$, we generate $\textbf{z}\in\mathbb{C}^{N}$ with random support with cardinality $n$ and the nonzero entries of z and $\textbf{H}\in\mathbb{C}^{k\times C}$ and $\textbf{z}\in\mathbb{C}^{N}$ yield $\mathcal{N}(0,1)$. We repeat the experiment ten times, and each time new H and z are generated.
We count an experiment as a success if the relative error between the solution of our optimization problem $\hat{\textbf{X}}$ and the truth $\textbf{X}_{0}$ is less than $1\%$, i.e., $\frac{\|\hat{\textbf{X}}-\textbf{X}\|_{F}}{\|\textbf{X}\|_{F}}\leq 1\%$. And the empirical success rate is a fraction defined as (total number of success) / (total number of experiments). We use the cvxpy [2][15] package of Python with the MOSEK [42] solver to solve the mixed norm optimization problem (17).
Even though in the theorems we treated B as a random matrix with row permutations, in the experiments we use a fixed B. We choose B as a fixed matrix consisting of the first $k$ columns of a random orthonormal matrix drawn from the $O(N)$ Haar distribution. It represents a stronger experiment than the theoretical treatment, as the theory only ensures recovery for most B, while experimentally, we show the performance of a more relevant circumstance in practice. We test the performance of our proposed method $\boldsymbol{\Psi}$ as the DCT matrix or wavelet basis. The results for the sparsifying basis as the DCT are shown in Figure 1.
Next, we illustrate the barrier for reconstruction when $\boldsymbol{\Psi}$ is the wavelet transform. From the theory, a small mutual coherence, i.e., $\mu^{\Psi}_{max}\sim O(1)$, is required for the performance of the compressive sensing method. However, the Daubechies wavelet, as the most widely used wavelet in sparsifying MRI images [13], has high coherence, indicating a barrier in the proposed method. We followed the non-uniform sampling scheme in the Fourier measurements as suggested by [13][1] for attempt to mitigate the problem. However, even when $L=N$ (when Fourier space is fully sampled), recovery is poor, as expected. We use the db4 wavelet as an illustration and the results are shown in Figure 2.
(a) 1 coil
(b) 2 coil
Figure 2:
5.1.1 Improved results compared to $l_{1}$ minimization
Our proposed method minimizes $\sum\limits_{i\in[N]}\|\textbf{X}_{S_{i}}\|_{2}$ to promote block sparsity within a lifted signal and between coils. We demonstrate its advantage over minimizing $\sum\limits_{i\in[C]}\|\textbf{X}(:,i)\|_{2}$, which promotes row sparsity between coils. We minimize $\sum\limits_{i\in[C]}\|\textbf{X}(:,i)\|_{2}$ in (17) and use the same experimental setup. The results are shown in Figure 3.
@@ 5.1.1 Improved results compared to $l_{1}$ minimization
Figure 3: Phase transition plot of performance by minimizing $\sum\limits_{i\in[C]}\|\textbf{X}(:,i)\|_{2}$. Observe that the improvement in the success rate is not obvious for including larger $C$. Compared to 1, the proposed optimization problem shows more improvement for larger $C$.
@@ 5.1.2 Minimal L required for exact recovery is proportional to $kn$
Next, we perform two sets of experiments to show that the minimal measurements $L$ required for exact recovery are proportional to $kn$.
  • For a fix $N=256$, $n=5$, we let $k$ ranges from 1 to 15 and $L$ varies from 10 to 200. We run the simulation for 10 times and calculate the empirical success rate
@@
We also repeat the same experiments for different $C=1,4$. The results are shown in the Figure 4.
@@ 5.2 Reconstruction of the analytical phantom
For the second part of the simulation, we focus on the MRI application. We point out two major differences from the last section:
  • As a crutial difference to the simulation, we do not manually enforce the image to be sparse/transform sparse so to allow for sparsity deficiency.
  • We account for the presence of noise.
@@
  • Instead of solving the constraint optimization problem (17), we solve the regularized problem (18).
@@
We use the analytical phantom of size $256$ by $256$ and randomly generated coil sensitivities following the Biot–Savart law. It is important to note that we do not manually enforce sparsity in our signal. For example, one may manually use a sparse estimation of the phantom x, e.g., by crafting the signal by thresholding $\textbf{z}=\boldsymbol{\Psi}\textbf{x}=0,if|\boldsymbol{\Psi}\textbf{x}|<\tau$ and set $\textbf{x}=\boldsymbol{\Psi}^{*}\textbf{z}$. The thresholding effectively reduces the dimensionality of the problem and would favor the reconstruction results. However, such manipulation is only possible in simulations when the true signal is known beforehand. Our approach is more flexible and allows for sparsity deficiency in the underlying signal.
@@
We use different number of coils to show the efficiency of the proposed method. We plot reconstruction results using different levels of sub-sampling rate and Gaussian noise level (i.e., noise to signal ratio). We report the relative error of the reconstructed image and the true phantom.
(a) 1 coil
@@
Figure 6: The x-axis and y-axis represents varying noise-to-signal ratios and the sub-sampling rate respectively. Our observation indicates a gradual degradation in reconstruction quality with increased noise and reduced sampling.
@@ 6 Proofs
6.1 Proof of Lemma (4)
Recall that $\boldsymbol{\Phi}=\begin{bmatrix}(\bf{\bar{q}}_{1}\bigotimes\bf{b}_{1})^{*}\\ \vdots\\ \bf{\bar{q}}_{N}\bigotimes\bf{b}_{N})^{*}\end{bmatrix}$, where each row of $\boldsymbol{\Phi}$ depends only on one row of B and $\boldsymbol{\Psi}^{*}$. Due to symmetry, the probability distribution of $\mathbb{\Phi}$ is unchanged if we assume B is fixed but the rows of $\mathbf{\Psi}^{*}$ are random. More precisely, we assume the rows of $\mathbf{\Psi}^{*}\in\mathbb{C}^{N\times M}$ are chosen uniformly at random without replacement from some $\mathbf{\Psi}_{0}^{*}\in\mathbb{C}^{N\times M}$ with $\mathbf{\Psi}_{0}^{*}\mathbf{\Psi}_{0}=\textbf{I}_{N}$. Under such a model, the rows of $\mathbf{\Psi}^{*}$ are not independent, making it difficult to analyze the probability model on $\boldsymbol{\Phi}$ directly. On the other hand, if the rows of $\boldsymbol{\Psi}$ are chosen uniformly at random with replacement from $\boldsymbol{\Psi}_{0}$, we can apply a matrix Bernstein inequality to estimate the operator norm of $\boldsymbol{\Psi}_{S}^{*}\boldsymbol{\Psi}_{S}-\textbf{I}_{S}$. We first state the following Lemma which follows almost exactly from Lemma 4.3 in [36]. The only minor change in the proof is in the estimation of the bounds of $R$ and $\delta^{2}$ for the matrix Bernstein inequality.
@@ 6.1 Proof of Lemma (4)
For any fixed matrix $\mathbb{B}\in\mathbb{C}^{N\times k}$ and $\boldsymbol{\Psi}^{*}\in\mathbb{C}^{N\times M}$ with its rows chosen uniformly at random with replacement from the rows of $\boldsymbol{\Psi}_{0}\in\mathbb{C}^{N\times M},\boldsymbol{\Psi}_{0}^{*}\boldsymbol{\Psi}_{0}=\mathbf{I}_{N}$. For the matrix $\boldsymbol{\Phi}$ defined in 3.1,
@@
The following Lemma relate Lemma (9) to the desired result.
@@ 6.2 Proof of Lemma (5)
where $\sigma=\|\sum\limits_{l=1}^{L}\mathbf{\mathbf{\Phi}}_{S}^{*}\textbf{a}_{l}\textbf{a}_{l}^{*}\mathbf{\mathbf{\Phi}}_{S}\mathbf{\mathbf{\Phi}}_{S}^{*}\textbf{a}_{l}\textbf{a}_{l}^{*}\mathbf{\mathbf{\Phi}}_{S}\|_{2\rightarrow 2}^{1/2}\leq\max\limits_{l}\{|\mathbf{\mathbf{\Phi}}_{S}^{*}\textbf{a}_{l}|\}\|\textbf{F}_{\Omega}\mathbf{\mathbf{\Phi}}_{S}\|_{2\rightarrow 2}=\|\mathbf{\mathbf{\Phi}}_{S}^{*}\textbf{F}_{\Omega}^{*}\|_{1\rightarrow 2}\|\textbf{F}_{\Omega}\mathbf{\mathbf{\Phi}}_{S}\|_{2\rightarrow 2}$. The derivation of the bound can be found, e.g. in Proposition 8.20 of [21].
By Proposition 7.13 in [21] regarding the moment estimation,
@@
Apply Lemma 14.3 [21] on the non-square matrix $\mathbf{\mathbf{\Phi}}_{S}^{*}\textbf{F}_{\Omega}^{*}$, for any $p\geq 1$,
@@
Then, we apply a tail bound from moment estimation, see Proposition 7.15 from [21]. For $u\geq 1$,
@@
This completes the proof for the Bernoulli case. For the case when the rows of $\textbf{F}_{\Omega}$ are chosen uniformly from $F$, use a similar argument as in the later part of the proof of Theorem 14.1 [21], where it shows the probability for the uniform model is bounded by twice the one for the Bernoulli model derived here.

### v2

@@ 4.4 Dual Certificate
Regarding the exact dual certificate, let $\textbf{V}=\textbf{A}_{S}(\textbf{A}_{S}^{*}\textbf{A}_{S})^{-1}\text{sgn}(\textbf{X}_{0})$, we have $\boldsymbol{\mathcal{Y}}=\textbf{A}^{*}\textbf{A}_{S}(\textbf{A}_{S}^{*}\textbf{A}_{S})^{-1}\text{sgn}(\mathbf{X}_{0})$ is the exact dual certificate. The next lemma deals with bounding the parameter $\tau$ in $\|\textbf{V}\|_{F}\leq\tau\sqrt{s}$ in Lemma 3. The bound obtained in the following lemma does not depend on the randomness of H or $sgn(\textbf{X}_{0})$.
@@
Let $\textbf{V}=\textbf{A}_{S}(\textbf{A}_{S}^{*}\textbf{A}_{S})^{-1}\text{sgn}(\textbf{X}_{0})$. Conditioned on $\|\textbf{A}_{S}^{*}\textbf{A}_{S}-\textbf{I}_{S}\|_{2\rightarrow 2}\leq\eta$,
@@
Remark: Note that increasing the number of channels $C$ increases the probability of recovery. Furthermore, the condition on $L$ is automatically met due to the bound $L\sim O(kn\log(k^{2}n^{2}N))$ from Lemma 5.
Equivalently, we can state the result in Lemma 8 as $\max\limits_{j\in[|S^{c}|]}\|\mathcal{P}_{S^{c}_{j}}\boldsymbol{\mathcal{Y}}\|_{F}<1/2$ with probability at least $1-N^{-\alpha}$ if
@@
5 Simulations
In the first part of the simulations, we demonstrate the sufficient conditions of $L$ and $C$ for exact recovery using simulated data with different underlying dimensions of $k$ and $n$. We calculate the empirical probability of success for different values of $k$ and $n$ and show that the transition curves for $L$ and $C$ are consistent with the analytically derived bounds. In the second part, we present reconstruction results using simulated data and real pMRI measurements.
@@ 5.1 Empirical success rate for the proposed method
(d) 8 coil
Figure 3: Repeated the same experimental setup as in Fig:2, with 1% of noise added.
For the first set of experiments, we follow the same setup as in Ling and Strohmer [ling_strohmer_2015] and calculate the empirical probability of success of the proposed method. Fix the number of measurements $L=128$ and the underlying signal dimension $N=256$, and solve the optimization problem (28) for varying $(k,n),k,n\in[15]$ and $C\in\{1,2,4\}$. For each $C$ and $(k,n)$, we generate $\textbf{z}\in\mathbb{C}^{N}$ with random support with cardinality $n$ and the nonzero entries of z and $\textbf{H}\in\mathbb{C}^{k\times C}$ drawn from $\mathcal{N}(0,1)$. Each time, we draw $L$ points uniformly from the Fourier space. We repeat the experiment ten times, and each time new H and z are generated.
We count an experiment as a success if the relative error between the solution of our optimization problem $\hat{\textbf{X}}$ and the truth $\textbf{X}_{0}$ is less than $0.01\%$, i.e., $\frac{\|\hat{\textbf{X}}-\textbf{X}_{0}\|_{F}}{\|\textbf{X}_{0}\|_{F}}\leq 10^{-4}$. And the empirical success rate is a fraction defined as (total number of successes) / (total number of experiments). We use the cvxpy [agrawal2018rewriting, diamond2016cvxpy] package of Python with the MOSEK [mosek] solver to solve the mixed norm optimization problem (28).
Even though in the theorems we treated B as a random matrix with row permutations, in the experiments we use a fixed B. We choose B as a fixed matrix consisting of the first $k$ columns of a random orthonormal matrix drawn from the $O(N)$ Haar distribution. It represents a stronger experiment than the theoretical treatment, as the theory only ensures recovery for most B, while experimentally, we show the performance of a more relevant circumstance in practice. The results for choosing the DCT matrix as the sparsifying transformation (i.e., $\boldsymbol{\Psi}$) are shown in Fig:2 and Fig:3 for the noise-free and noisy cases, respectively.
Next, we illustrate the barrier for reconstruction with the proposed framework when $\boldsymbol{\Psi}$ is a Wavelet transform. From the theory, a small mutual coherence, i.e., $\mu^{\Psi}_{max}\sim O(1)$, is required for the performance of the compressive sensing method. However, the Daubechies wavelet, which can be used to sparsify MRI images [Chun_Adcock_Talavage_2014], has high coherence. We followed the non-uniform sampling scheme in the Fourier measurements as suggested by [Chun_Adcock_Talavage_2014, ADCOCK_HANSEN_POON_ROMAN_2017] as an attempt to improve the results. However, even when $L=N$ (when Fourier space is fully sampled), recovery is poor. We use the Daubechies 4 wavelet as an illustration, and the results are shown in Figure 4.
(a) 1 coil
(b) 2 coil
Figure 4: Phase transition plot of performance by solving (28) and using db4 wavelet transform as the sparsifying transformation matrix.
Nevertheless, since our method is designed to work with approximately sparse signals, and given that we will demonstrate the effectiveness of using the Discrete Cosine Transform (DCT) as the sparsifying transformation, it is not critical if the wavelet does not perform as well in this context.
5.1.1 Improved results compared to $\ell_{1}$ minimization
Our proposed method minimizes $\sum\limits_{i\in[N]}\|\textbf{X}_{S_{i}}\|_{2}$ to promote block sparsity within a lifted signal and between coils. We demonstrate its advantage over minimizing $\sum\limits_{i\in[C]}\|\textbf{X}(:,i)\|_{2}$, which only promotes row sparsity between coils. We change the objective in (28) to minimize $\sum\limits_{i\in[C]}\|\textbf{X}(:,i)\|_{2}$ and use the same experimental setup. The results are shown in Figure 5.
(a) 1 coil
(b) 2 coil
(c) 4 coil
Figure 5: Phase transition plot of performance by minimizing $\sum\limits_{i\in[C]}\|\textbf{X}(:,i)\|_{2}$ instead. Observe that compared to the proposed method, the improvement in the success rate is not obvious for including larger $C$.
@@ 5.1.2 Minimal L required for exact recovery is proportional to $kn$
Next, we perform two sets of experiments to show that the minimal measurements $L$ required for exact recovery are proportional to the sparsity level $kn$.
  • For a fix $N=256$, $n=5$, we let $k$ ranges from 1 to 15 and $L$ varies from 10 to 200. We run the simulation for 10 times and calculate the empirical success rate.
@@
We also repeat the same experiments for different $C=1,4$. The results are shown in Figure 6 and 7.
@@ 5.2 Reconstruction of the analytical phantom
For the second part of the simulation, we show reconstruction results using the analytical phantom and simulated coils. We point out the major differences of the experimental setup from the last part:
  • A crucial difference in the simulations is that we do not manually enforce the object of interest (i.e. x or z) to be sparse/transform sparse. Hence, it allows for sparsity deficiency.
  • Different levels of noise are added to the measurements.
@@
  • Instead of solving the constrained optimization problem (28), we solve the regularized problem (29) using an ISTA-type of updates as shown in Appendix 7.2.
@@
We use the analytical phantom [Guerquin-Kern_Lejeune_Pruessmann_Unser_2012] of size $256$ by $256$ and randomly generated coil sensitivities following the Biot–Savart law. It is important to note that we do not manually enforce sparsity in our signal. For example, one may manually use a sparse estimation of the phantom x, e.g., by crafting the signal by thresholding $\textbf{z}=\boldsymbol{\Psi}\textbf{x}=0,if|\boldsymbol{\Psi}\textbf{x}|<\tau$ and set $\textbf{x}=\boldsymbol{\Psi}^{*}\textbf{z}$. The thresholding effectively reduces the dimensionality of the problem and would favor the reconstruction results. However, such manipulation is only possible in simulations when the true signal is known beforehand. Our approach is more flexible and allows for sparsity deficiency in the underlying signal.
@@
We use different number of coils to show the efficiency of the proposed method. We plot reconstruction results using different levels of sub-sampling rate and Gaussian noise level (i.e., noise to signal ratio). We report the relative error (rerr) of the reconstructed image and the true phantom. The reconstruction results are shown in Figure 9.
(a) 2 coils
(b) 4 coils
(a) 8 coils
@@
Figure 9: Reconstruction of the analytical phantom. The x-axis and y-axis represents varying noise-to-signal ratios and the sub-sampling rate respectively. Our observation indicates a gradual degradation in reconstruction quality with increased noise and increased reduction factor.
To assess the performance of the proposed method with respect to the number of coils, we plot the average reconstruction error versus the number of measurements per coil, for different numbers of coils and levels of Gaussian noise. We adopted a similar setup as that in [otazo2009distributed, Otazo_Kim_Axel_Sodickson_2010] where the multi-coil signal was generated by multiplying a small analytical phantom (N = 529) by the coil sensitivities (k=6) and optionally adding Gaussian noise. The average reconstruction error (relative $\ell_{2}$ distance) over 10 different random Fourier space undersampling realizations was computed for different numbers of measurements and coil elements. A DCT transform was employed to sparsify the image and no strict sparsity was enforced in the signal. The number of required measurements to achieve the noise floor reconstruction decreases with the number of coils for both the noise-free and the noisy case. These trends are consistent with previous numerical experiments for the known calibration pMRI reconstruction, e.g. [otazo2009distributed]. The plots confirm that with a large number of coils, the number of measurements per coil sufficient to achieve a baseline level of reconstruction error is significantly smaller than the true dimensionality of the signal.
Figure 10: The four subplots correspond to the noise-free case and SNR values of 100, 50, and 10. The relative $\ell_{2}$ error of the reconstruction is plotted against varying number of measurements per coil ($L$). For SNR=10, the log scale of the relative error is plotted in the inset of (d).
5.3 Comparison with the ENLIVE method
Next, we compare our reconstruction results with the ENLIVE formulation (17) in the paper by Holme et al. [Holme_Rosenzweig_Ong_Wilke_Lustig_Uecker_2019]. We reproduced the reconstructed brain and knee images exactly as shown in the original paper using the data and source code provided in https://github.com/mrirecon/enlive. We then compared these results with our own reconstructions. Figures 11 and 12 show calibrationless variable-density Poisson-disc undersampled reconstructions with differing undersampling factors comparing ENLIVE to our proposed method. Figure 13 shows the reconstructions of the simulated analytical phantom from uniformly subsampled Fourier space with different undersampling factors. As a reference, the corresponding sampling patterns are shown in the third columns of each of the plots.
Figure 11: Variable-density Poisson-disc undersampled data with varying undersampling factors were reconstructed using ENLIVE (with $k=2$ in (17)) as well as our proposed method. For undersampling factors up to $R=7.0$, both methods produce images that are essentially realistic. However, at an undersampling factor of $R=8.5$, the ENLIVE reconstruction becomes very noisy, and creates some features void in the center of the image. With higher undersampling factors, artifacts appear in the images reconstructed with our proposed method, but the brain retains more of its features and shows a more uniform contrast.
Figure 12: Variable-density Poisson-disc undersampled data of a human knee with varying undersampling factors reconstructed with ENLIVE allowing two sets of maps (i.e. with $k=2$ in eq:(17)) and with our method. The proposed method provides reconstructions with higher and more consistent contrast. For R = 5.0, the reconstruction by ENLIVE seems preferable.
Figure 13: Uniformly undersampled noisy data of the analytical phantom and simulated coils with varying undersampling factors reconstructed with ENLIVE and our method. Relative $\ell_{2}$ error of the reconstructed phantom and the true phantom is computed. For reduction factors up to $R=5.0$, the ENLIVE produces reconstructions with low signal at the center, whereas our method produces images with uniform contrast throughout. The reconstructions of our method all have smaller $\ell_{2}$ error than the ENLIVE reconstructions. For $R=5.0$, artifacts start to appear in our reconstruction.
@@ 6 Proofs
6.1 Proof of Lemma 3
Recall that $\boldsymbol{\Phi}=\begin{bmatrix}(\bf{\bar{q}}_{1}\bigotimes\bf{b}_{1})^{*}\\ \vdots\\ \bf{\bar{q}}_{N}\bigotimes\mathbf{b}_{N})^{*}\end{bmatrix}$, where each row of $\boldsymbol{\Phi}$ depends only on one row of B and $\boldsymbol{\Psi}^{*}$. Due to symmetry, the probability distribution of $\mathbb{\Phi}$ is unchanged if we assume B is fixed but the rows of $\mathbf{\Psi}^{*}$ are random. More precisely, we assume the rows of $\mathbf{\Psi}^{*}\in\mathbb{C}^{N\times M}$ are chosen uniformly at random without replacement from some $\mathbf{\Psi}_{0}^{*}\in\mathbb{C}^{N\times M}$ with $\mathbf{\Psi}_{0}^{*}\mathbf{\Psi}_{0}=\textbf{I}_{N}$. Under such a model, the rows of $\mathbf{\Psi}^{*}$ are not independent, making it difficult to analyze the probability model on $\boldsymbol{\Phi}$ directly. On the other hand, if the rows of $\boldsymbol{\Psi}$ are chosen uniformly at random with replacement from $\boldsymbol{\Psi}_{0}$, we can apply a matrix Bernstein inequality to estimate the operator norm of $\boldsymbol{\Psi}_{S}^{*}\boldsymbol{\Psi}_{S}-\textbf{I}_{S}$. We first state the following Lemma which follows almost exactly from Lemma 4.3 in [ling_strohmer_2015]. The only minor change in the proof is in the estimation of the bounds of $R$ and $\delta^{2}$ for the matrix Bernstein inequality.
@@ 6.1 Proof of Lemma 3
For any fixed matrix $\mathbb{B}\in\mathbb{C}^{N\times k}$ and $\boldsymbol{\Psi}^{*}\in\mathbb{C}^{N\times M}$ with its rows chosen uniformly at random with replacement from the rows of $\boldsymbol{\Psi}_{0}\in\mathbb{C}^{N\times M},\boldsymbol{\Psi}_{0}^{*}\boldsymbol{\Psi}_{0}=\mathbf{I}_{N}$. For the matrix $\boldsymbol{\Phi}$ defined in (24),
@@
The following lemma relates Lemma 9 to the desired result.
@@ 6.2 Proof of Lemma 4
where $\sigma=\|\sum\limits_{l=1}^{L}\mathbf{\mathbf{\Phi}}_{S}^{*}\textbf{a}_{l}\textbf{a}_{l}^{*}\mathbf{\mathbf{\Phi}}_{S}\mathbf{\mathbf{\Phi}}_{S}^{*}\textbf{a}_{l}\textbf{a}_{l}^{*}\mathbf{\mathbf{\Phi}}_{S}\|_{2\rightarrow 2}^{1/2}\leq\max\limits_{l}\{|\mathbf{\mathbf{\Phi}}_{S}^{*}\textbf{a}_{l}|\}\|\textbf{F}_{\Omega}\mathbf{\mathbf{\Phi}}_{S}\|_{2\rightarrow 2}=\|\mathbf{\mathbf{\Phi}}_{S}^{*}\textbf{F}_{\Omega}^{*}\|_{1\rightarrow 2}\|\textbf{F}_{\Omega}\mathbf{\mathbf{\Phi}}_{S}\|_{2\rightarrow 2}$. The derivation of the bound can be found, e.g. in Proposition 8.20 of [foucart_rauhut_2015].
By Proposition 7.13 in [foucart_rauhut_2015] regarding the moment estimation,
@@
Apply Lemma 14.3 [foucart_rauhut_2015] on the non-square matrix $\mathbf{\mathbf{\Phi}}_{S}^{*}\textbf{F}_{\Omega}^{*}$, for any $p\geq 1$,
@@
Then, we apply a tail bound from moment estimation, see Proposition 7.15 from [foucart_rauhut_2015]. For $u\geq 1$,
@@
This completes the proof for the Bernoulli case. For the case when the rows of $\textbf{F}_{\Omega}$ are chosen uniformly from $F$, use a similar argument as in the later part of the proof of Theorem 14.1 [foucart_rauhut_2015], where it shows the probability for the uniform model is bounded by twice the one for the Bernoulli model derived here.